OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues.1 - 3 + 9 - 27 + ...

OpenStudy (anonymous):

the pattern is \((-1)^n3^n\)

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}1*3^n\] \[\sum_{n=0}^{\infty} 1*3^(n+1)\] \[\sum_{n=0}^{\infty} 1(-3)^n \] \[\sum_{n=0}^{\infty} 1(-3)^(n+1)\]

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

More like \[\sum (-1)^n3^n \]

OpenStudy (anonymous):

what about this one? Write the sum using summation notation, assuming the suggested pattern continues. 100 + 121 + 144 + 169 + ... + n2 + ...